Show that

∫ₐᵇ (ln x)/(x) dx = (1/2) ln ab ln(a/b)
1
Expert's answer
2019-11-01T12:21:59-0400

"\intop ^b_a \frac {ln(x)} x dx"

Find

"\int \frac {ln(x)} x dx = \int ln(x) d(ln(x))=[ln(x)=t]=\int t dt = \frac {t^2} 2 + C = \frac {ln^2(x)} 2 + C"

than

"\intop ^b_a \frac {ln(x)} x dx = \frac {ln^2(x)} 2|^b _a = \frac {ln^2(b)} 2 - \frac {ln^2(a)} 2 = \frac 1 2 (ln^2 (b) - ln^2(a)) = \frac 1 2 (ln(b) - ln(a))(ln(b) +ln(a)) = \frac 1 2 (ln (\frac b a)ln(ab))"